If you earn an annual interest rate of 10.3 percent, how many years will it take to quadruple your money?

14.14 years

13.05 years

12.86 years

11.31 years

12.37 years
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value($4x)
P=present value($x say)
r=rate of interest
n=time period.

4x=x*(1.103)^n

4=(1.103)^n

Taking log on both sides;

log 4=n*log (1.103)

n=log 4/log (1.103)

=14.14 years(Approx).